OpenStudy (anonymous):

A cube has a surface area of 96 square centimeters. How long is each face of the cube?

OpenStudy (anonymous):

A=6a^2 6x(?)=96 96/6 = 16 \[\sqrt{16}\] a=4

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

anyone else for a reply on my question?

OpenStudy (kaptain_mittens):

The formula for finding surface area of a cube is A=6a^2 If you need more help from here then ask but you should be able to solve from here.

OpenStudy (kaptain_mittens):

Oh too late, I was beat to the punch :)

OpenStudy (kaptain_mittens):

Can anyone try and solve my question now, I really need help

OpenStudy (bradely):

Surface of the cube s = 6a^2 6a^2=96 a^2=16 a=4 Source: http://www.mathskey.com/question2answer/ask

OpenStudy (anonymous):

thanks